Search Archive:
<< Previous      Next >>
804 Views 10 August Derecho Outage


Nothing short of a crazy day today. I'll update this news item frequently as I make progress restoring the IEM and filling the data void left today due to the Ames/ISU power outage from the Derecho of 2020!

11:21 PM: Power is back and services are getting back on their feet. Lots of issues yet and data holes to fill.

Updated 11 Aug 12:04 AM: The IEM should be functioning again for realtime requests. The AFOS (text product) database hole should now be filled. This is the first step in repairing other datasets on this genre. The extent of the outage was from about 11 AM this morning until 10 PM this evening.

Updated 11 Aug 12:22 AM: Some ISU ITS managed virtual machines remain offline and there is cooling capacity issues in the central data center. This marginally impacts the IEM, but some archiving is not working at the moment due to it.

Updated 5:30 AM: No power at my house is creating for a slow go with getting more of the data holes repaired. Will do what I can and keep this news item updating.

Updated 10:50 AM: Maybe I now have the hole in the VTEC (NWS Watch Warning Advisory) database and Local Storm Report (LSR) database filled. It is difficult to tell with my poor network working conditions at the moment.

Updated 8:45 PM: The logistical nightmare continues and repairs to the holes is slow going. Perhaps I have the RIDGE imagery backfilled now and will soon have the NEXRAD composites repaired. It is hopeless for me at the moment to catch up on emails and twitter mentions. Thanks for your patience.

Updated 12 August 4:30 AM: I believe the hole with the ASOS/METAR data should now be filled. I am getting questions about peak gusts in Iowa for the event and I can not really answer them until I get these holes repaired. Filling the RWIS one may not be possible as I am unsure if the DOT, who is located in Ames, was able to collect data during the Ames power outage. I will reach out to them soon.

Updated 8:05 AM: Repaired NEXRAD storm attribute archive hole.

Updated 10:35 PM: Not much for updates to share. I continue to struggle with logistics due to no power and Internet at home. Thankfully, power and cooling at work have been stable.

Update 13 Aug 12:00 PM: I have power and internet at home now, so can get after this task some more!

Update 13 Aug 4:00 PM: A previous backfill of VTEC / Storm Based Warning products did not go as planned. This was cleaned up and should be all square now.

Update 16 August 2:20 PM: I have now repaired the hole with the SPC Day 1 outlooks and MCDs.